I have to use (1quart =0.946)and convert 425ml to quarts I did 425ml x0.946L/1ml x 1quarts/0.946L..did I set this up right?...I got 425.000 I have to have the correct significant figures
1
Expert's answer
2015-09-11T03:08:40-0400
1 quart=0.946353 l=946.353 ml
1 ml = 1/946.353 quart
So, 425 ml = 425/946.353 = 0.449092 quart